OpenStudy (anonymous):

Solve the formula for the variable h. V=bh/3 i need help fast!!

OpenStudy (darkbluechocobo):

o_o are you trying to find h?

OpenStudy (anonymous):

no i just need to know the fourmula

OpenStudy (darkbluechocobo):

so you have to divide V by h so V/h=b/3

OpenStudy (anonymous):

would that = h = v/3b

OpenStudy (darkbluechocobo):

Yes you beat me to it xD

OpenStudy (darkbluechocobo):

let me check it quick though

OpenStudy (anonymous):

eally thats it! awesome

OpenStudy (darkbluechocobo):

Yes :p it is

OpenStudy (darkbluechocobo):

It has been nice working with you again xD

OpenStudy (anonymous):

yes thank you so much. fans?

OpenStudy (darkbluechocobo):

sures :p